RC Meeting, Sunday, October 24 , 2021

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920

Accelerator Status

  1. The accelerator still has an unresolved issue regarding A beam going thru the C slit and causing BLM trips in Arc 7. The problem was apparently dodged last nite by restricting Hall A to lower beam currents, but that won't last.
  2. Regarding the Arc 7 magnet short, on Monday there will be additional inspection and measurements for construction of a drift pipe IF the magnet is going to be pulled out. So after we recover beam on Monday, expect continued delivery to A,B,C but no Hall D.
  3. Pass change for Hall A is definitely Monday.
  4. Regarding the Secretary of Energy tour, VIP schedules can change unexpectedly, but the last info we were given was that the accelerator will be off 09:30 Monday morning. Tours wrap up at 13:30.

Hall C Status

  1. We had good production at ~42muA starting 18:15 on Saturday, and continuing into Sunday day shift.
  2. Immediately before that, there was no beam delivery to Hall C during 16:30-18:15. And most of Saturday day shift I would call QBU (questionable beam in use during 09:00-16:30) since our beam current varied from 6-40 muA while MCC was working the issue of BLM trips when trying to deliver requested currents to both Halls A and C.
  3. Number of good pions was 4629 as of 08:00 Sunday. See Ali's hclog https://logbooks.jlab.org/entry/3931371

Current Run Plan

  1. Continue data taking until early Nov

Current Issues

  1. Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
  2. Q3 insulating vacuum spikes generating alarms. For example, https://logbooks.jlab.org/entry/3930175 .
  3. One HB LHe incident of note: https://logbooks.jlab.org/entry/3929152 . (possible clue to the origin?)
  4. Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
  5. Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
  6. HCCAM02 not functioning (lift required to replace).

Special Requests

Useful Links